1. A jogger runs 224 meters eastward, then 156 meters westward, then 84 meters eastward, and finally 248 meters westward.

For this motion, what is the distance moved? _________________

What is the magnitude and direction of the displacement? _________________

2. A bowling ball moves 15 meters northward, 12 meters southward, then 27 meters northward, and finally 5 meters southward.

For this motion, what is the distance moved? _________________

What is the magnitude and direction of the displacement? _________________


1
Expert's answer
2021-11-14T17:18:04-0500

1) For this motion, what is the distance moved? 224+156+84+248=712 m.

What is the magnitude and direction of the displacement? 224-156+84-248=-96 m E, or 96 m to the West.


2) For this motion, what is the distance moved? 15+12+27+5=59 m

What is the magnitude and direction of the displacement? 15-12+27-5=25 m northward.